# Learning Neural Control Barrier Functions from Offline Data with Conservatism

### v1

@@ B.2 Learning a NCBF from data labeled using an extremely conservative strategy
The empirical findings indicate that, for both model sizes, on average across all nominal controllers, this extreme labeling method results in poorer performance compared to the nominal policy without any safety filter. Notably, across both Swimmer and Hopper environments, pairing BEAR-L with this extreme version of iDBF achieves lower cost (with similar reward levels) than using BEAR-L without the safety filter. However, combining COptiDICE or BCQ-L with iDBF severely degrades performance, as evidenced by a significant increase in cost. We note that these findings align with those in [60] which suggest that assigning high-cost values to both OOD and unsafe actions can result in poor generalizability, which might explain the suboptimal performance observed during policy roll-outs when training a safety filter with this labeling strategy.
@@
Appendix C Additional discussion of results: safety barrier performance is dependent on the nominal controller and the environment
However, in the vision-based navigation experiment, the choice of nominal controller had a pronounced impact on performance. We can see in Table 1 how all the learned barrier functions perform poorly when paired with the goal reaching PD nominal controller . However, when paired with BC and BC-Safe nominal controllers, both iDBF and CCBF perform well, with CCBF achieving zero collisions and high success rate.
Appendix D Visual representation of the learned CCBF and iDBF using different hyperparameters in the 2D navigation environment
(a) Comparison of different CCBFs learned with different values of $w_{c}$ shows that incorporating $\mathcal{L}_{c}$ makes the CCBF more conservative than NCBF, while still preserving its ability to distinguish between safe and unsafe states. When $w_{c}=1$, the CCBF effectively differentiates between InD and OOD states.
(b) Comparison of different iDBFs learned with different density thresholds $p$ to select out-of-distribution actions highlights that iDBF is sensitive to the choice of $p$, and struggles to strike a proper balance between accurately separating safe and unsafe sets while respecting the underlying training distribution.
Figure 4: Visual comparison of learned CCBF and iDBF models with different hyperparameters.
Appendix E Closed loop simulation results of CCBFs learned with and without detaching $B_{\phi}(x^{\prime})$ in the computation of $L_{c}$ during training
Table 10: Evaluation results for different pairs of nominal controllers and CCBFs trained with and without detaching $B_{\phi}(x^{\prime})$ when computing $L_{c}$ during training. Results of each pair are averaged across 500 runs for the 2D navigation task and 50 runs for the vision-based navigation task.
Policy None NCBF (without detach) CCBF (with detach)
Success % Collision % Success % Collision % Success % Collision %
2D Navigation
PD controller 91.6±3.3 0 91.0±2.2 0
BC 96.2±2.2 21.8±4.2 96.4±2.3 0 95.4±2.6 0
BC‑Safe 93.6±3.2 6.4±2.3 93.2±3.5 0 92.8±3.6 0
Average 94.9 14.1 93.7 0 93.1 0
Vision‑based Navigation
PD controller 68.0±17.9 82.0±8.4 58.0±14.8 4.0±5.5
BC 98.0±4.5 36.0±16.7 58.0±4.5 34.0±16.7 86.0±11.4 0
BC‑Safe 92.0±8.4 22.0±4.5 62.0±4.5 24.0±8.9 94.0±8.9 0
Average 95.0 29.0 62.7 46.7 79.3 1.3
@@
Policy None CCBF (without detach) CCBF (with detach)
@@
BC 2.26±0.64 0.44±0.03 0.61±0.17 0.31±0.02 0.94±0.18 0.39±0.03
BC-Safe 0.12±0.05 0.43±0.03 0.07±0.01 0.38±0.01 0.03±0.02 0.45±0.02
COptiDICE 1.65±0.33 0.30±0.04 0.05±0.04 0.39±0.04 0.19±0.08 0.28±0.06
BEAR-L 0.61±0.11 0.16±0.03 0.48±0.09 0.13±0.02 0.42±0.17 0.09±0.01
BCQ-L 2.82±1.15 0.25±0.08 2.91±1.31 0.32±0.03 1.40±0.47 0.29±0.03
Average 1.49 0.32 0.82 0.31 0.59 0.30
@@
BC 0.19±0.27 0.04±0.02 0.09±0.05 0.04±0.01 0.09±0.06 0.04±0.01
BC-Safe 0.03±0.02 0.57±0.01 0.16±0.11 0.58±0.02 0.05±0.03 0.56±0.05
COptiDICE 0.01±0.01 0.18±0.01 0.01±0.01 0.17±0.01 0.03±0.03 0.17±0.01
BEAR-L 0.37±0.01 0.16±0.01 0.18±0.01 0.15±0.01 0.20±0.06 0.30±0.06
BCQ-L 3.09±0.34 0.50±0.04 1.43±0.15 0.31±0.04 1.16±0.53 0.37±0.05
Average 0.74 0.29 0.37 0.25 0.31 0.28
It can be seen that detaching $B_{\phi}(x^{\prime})$ in the computation of $L_{c}$ during training of CCBFs leads to better performance compared to when not detaching it. In the 2D navigation task, both CCBFs (with and without detach) perform similarly, whereas CCBF (with detach) performs better in the vision-based navigation task. In both Swimmer and Hopper environments, CCBF (without detach) underperforms CCBF (with detach), but is still able to significantly improve the closed loop performance of the nominal controllers and outperforms NCBF and iDBF.

### v2

@@ A-B Learning iDBF with density threshold $p=1$
The empirical findings indicate that, for both model sizes, on average across all nominal controllers, this extreme labeling method results in poorer performance compared to the nominal policy without any safety filter. Notably, across both Swimmer and Hopper environments, pairing BEAR-L with this extreme version of iDBF achieves lower cost (with similar reward levels) than using BEAR-L without the safety filter. However, combining COptiDICE or BCQ-L with iDBF severely degrades performance, as evidenced by a significant increase in cost.
@@
A-C Closed loop evaluation of CCBF* paired with different controllers
we studied whether other conservativeness-inducing loss terms would result in similar improvement in performance compared to the NCBF approach. We denote the first candidate loss term, which replaces $\mathcal{L}_{c}$, by $\mathcal{L}_{\text{unsafe}}^{\prime}$ and its associate weight by $w_{\text{unsafe}}^{\prime}$. The term $\mathcal{L}_{\text{unsafe}}^{\prime}$ is equivalent $\mathcal{L}_{\text{unsafe}}$ evaluated at the states sampled in the same manner when computing $\mathcal{L}_{c}$, i.e., treating the states reached by randomly-sampled actions similarly to unsafe ones. We call this new approach CCBF.
@@ A-C Closed loop evaluation of CCBF* paired with different controllers
Policy None CCBF* ($w_{c}^{*}=0.5$) CCBF* ($w_{c}^{*}=1$)
@@
BC 2.26±0.64 0.44±0.03 2.06±0.16 0.37±0.03 2.38±0.66 0.40±0.04
BC‑Safe 0.12±0.05 0.43±0.03 0.13±0.08 0.35±0.01 0.06±0.02 0.43±0.02
COptiDICE 1.65±0.33 0.30±0.04 0.96±0.42 0.48±0.01 6.01±0.82 0.35±0.03
BEAR‑L 0.61±0.11 0.16±0.03 0.46±0.09 0.15±0.02 0.56±0.16 0.17±0.02
BCQ‑L 2.82±1.15 0.25±0.08 3.03±0.50 0.33±0.02 3.20±0.91 0.26±0.04
Average 1.49 0.32 1.33 0.34 2.44 0.32
@@
BC 0.19±0.27 0.04±0.02 0.39±0.09 0.07±0.01 1.71±0.33 0.18±0.02
BC‑Safe 0.03±0.02 0.57±0.01 0.11±0.06 0.56±0.01 0.27±0.16 0.52±0.02
COptiDICE 0.01±0.01 0.18±0.01 0.08±0.03 0.18±0.01 0.16±0.05 0.18±0.01
BEAR‑L 0.37±0.01 0.16±0.01 0.37±0.17 0.39±0.01 0.24±0.13 0.26±0.00
BCQ‑L 3.09±0.34 0.50±0.04 0.92±0.28 0.34±0.01 2.15±05 0.18±0.04
Average 0.74 0.29 0.37 0.31 0.91 0.26
A-D Additional discussion of results: safety barrier performance is dependent on the nominal controller and the environment
However, in the vision-based navigation experiment, the choice of nominal controller had a pronounced impact on performance. We can see in Table I how all the learned barrier functions perform poorly when paired with the goal reaching PD nominal controller . However, when paired with BC and BC-Safe nominal controllers, both iDBF and CCBF perform well, with CCBF achieving zero collisions and high success rate.
A-E Comparison of the learned neural CBFs using CCBF, CCBF*, and NCBF in the Hopper, Swimmer, and vision-based navigation task
We evaluate and compare the neural control barrier functions learned using NCBF, CCBF, and CCBF* across the Hopper, Swimmer, and vision-based navigation tasks. For each method, we examine how the barrier function values differ between safe and unsafe states in test trajectories, as well as states reached by taking either dataset actions or randomly sampled actions starting from safe states in the test trajectories.
Appendix B Comparison of the learned neural CBFs using different methods in the Hopper task
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Figure 3: Variation of the mean neural control barrier function values for NCBF, CCBF, and CCBF* on Hopper test trajectories. Left column: safe vs. unsafe states in test trajectories. Right column: states reached by taking either dataset actions or randomly sampled actions from safe states in unseen test trajectories.
Appendix C Comparison of the learned neural CBFs using different methods in the Swimmer task
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Figure 4: Variation of the mean neural control barrier function values for NCBF, CCBF, and CCBF* on Swimmer test trajectories. Left column: safe vs. unsafe states in test trajectories. Right column: states reached by taking either dataset actions or randomly sampled actions from safe states in unseen test trajectories.
Appendix D Comparison of the learned neural CBFs using different methods in the vision-based navigation task
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Figure 5: Variation of the mean neural control barrier function values for NCBF and CCBF on vision-based navigation test trajectories. Left column: safe vs. unsafe states in test trajectories. Right column: states reached by taking either dataset actions or randomly sampled actions from safe states in unseen test trajectories.